### v1

# A note on Lie and Jordan structures of Leavitt path algebras

Abstract.
Let $L_{K}(E)$ be the Leavitt path algebra of a directed graph $E$ over a field $K$. In this paper, we determine $E$ and $K$ for the Lie algebra $\mathbf{K}_{L_{K}(E)}$ and the Jordan algebra $\mathbf{S}_{L_{K}(E)}$ arising from $L_{K}(E)$ with respect to the standard involution to be solvable.
Key words and phrases:
leavitt path algebra; cohn path algebra; Lie algebra; Jordan algebra
2020 Mathematics Subject Classification. 16S88; 16W10
Corresponding author
1. Introduction
Every associative algebra $\mathcal{A}$ over a field $K$ can give rise to two new structures: the Lie algebra, denoted by $\mathcal{A}^{-}$, and the Jordan algebra, denoted by $\mathcal{A}^{\circ}$ over $K$. These structures can be obtained by defining new products within $\mathcal{A}$, namely the Lie bracket $[x,y]=xy-yx$ for Lie algebras and the Jordan product $x\circ y=xy+yx$ for Jordan algebras respectively, for every $x,y\in\mathcal{A}$. A reasonable question arises: Are the properties of $\mathcal{A}$ as a ring mirrored in its Lie or Jordan algebra counterparts? Specifically, does the simplicity of $\mathcal{A}$ as a ring guarantee its simplicity as a Lie or Jordan algebra? In a seminal paper [9] by Herstein, this question was answered affirmatively. Assume further that $\mathcal{A}$ is equipped with an involution $\star$; that is, a map ${}^{\star}:\mathcal{A}\to\mathcal{A}$ satisfying all below conditions:
  1. (1)
    $(a+b)^{\star}=a^{\star}+b^{\star}$,
  2. (2)
    $(ab)^{\star}=b^{\star}a^{\star}$,
  3. (3)
    $(a^{\star})^{\star}=a$,
for all $a,b\in\mathcal{A}$. Let
$\mathbf{S}_{\mathcal{A}}=\{x\in\mathcal{A}:x^{\star}=x\}\text{ and }\mathbf{K}_{\mathcal{A}}=\{x\in\mathcal{A}:x^{\star}=-x\}$
[…]
In this paper, we consider the case when $\mathcal{A}$ is the Leavitt path algebra $L_{K}(E)$ of a directed graph $E$ over a field $K$. Within the past few years, the Lie subalgebras of $L_{K}(E)^{-}$ arising from $L_{K}(E)$ have been investigated by many authors. Motivated by the groundbreaking work of Herstein [9], necessary and sufficient conditions for the Lie algebra $[L_{K}(E),L_{K}(E)]$ of $L_{K}(E)^{-}$ to be simple were given in [2] and [3], under the assumption that $E$ is row-finite. In another direction, the paper [12] completely identified the graph $E$ and the field $K$ for the Lie algebra $L_{K}(E)^{-}$ to be solvable.
Let $E$ be a graph and $K$ a field. Then, the mapping ${}^{\star}:L_{K}(E)\to L_{K}(E)$ which sends $v$ to $v$, $e$ to $e^{*}$ and $e^{*}$ to $e$ for all $v\in E^{0}$ and $e\in E^{1}$, and leaves all elements in $K$ fixed, extends to an involution of $L_{K}(E)$. We call this the standard involution on $L_{K}(E)$ (see also [1, page 36]). Let $\mathbf{K}_{L_{K}(E)}$ and $\mathbf{S}_{L_{K}(E)}$ be the Lie subalgebra of $L_{K}(E)^{-}$ and the Jordan subalgebra of $L_{K}(E)^{\circ}$ with respect to this involution, respectively. In [4] the simplicity of the Lie algebra $\mathbf{K}_{L_{K}(E)}$ was also studied. The current paper is devoted to a study of the Lie algebra $\mathbf{K}_{L_{K}(E)}$ and the Jordan algebra $\mathbf{S}_{L_{K}(E)}$. To be more specific, we identify the graph $E$ and the field $K$ for the Lie algebra $\mathbf{K}_{L_{K}(E)}$ and the Jordan algebra $\mathbf{S}_{L_{K}(E)}$ to be solvable.
2. Preliminaries
2.1. Directed graphs and Leavitt path algebras.
Throughout this paper, the basic notations and conventions are taken from [1]. In particular, a (directed) graph $E=(E^{0},E^{1},r,s)$ consists of two sets $E^{0}$ and $E^{1}$ together with maps $r,s:E^{1}\to E^{0}$. The elements of $E^{0}$ are called vertices and the elements of $E^{1}$ edges. For the convenience of readers, in the following, we outline some concepts and notations that we use substantially in this paper.
Let $E$ be a graph. A vertex $v$ in $E$ is a sink if it emits no edges, while it is an infinite emitter if it emits infinitely many edges. A graph $E$ is said to be row-finite if $E$ contains no infinite emitter. A vertex $v$ is said to be regular if it is neither a sink nor an infinite emitter. A path $\mu$ in a graph $E$ is a finite sequence of edges $\mu=e_{1}e_{2}\dots e_{n}$ with $r(e_{i})=s(e_{i+1})$ for all $11\leq i\leq n-$ 1. We set $s(\mu):=s(e_{1})$ and $r(\mu):=r(e_{n})$. The set of all finite paths in $E$ is denoted by ${\rm Path}(E)$. An exit for a path $\mu=e_{1}\dots e_{n}$ is an edge $e$ such that $s(e)=s(e_{i})$ for some $i$ and $e\neq e_{i}$.
Definition 2.1 (Leavitt path algebra).
Let $E$ be a graph and $K$ a field. We define a set $\left(E^{1}\right)^{*}$ consisting of symbols of the form $\{e^{*}|e\in E^{1}\}$. The Leavitt path algebra of $E$ with coefficients in $K$, denoted by $L_{K}(E)$, is the free associative $K$-algebra generated by $E^{0}\cup E^{1}\cup(E^{1})^{*}$, subject to the following relations:
  1. (V)
    $vv^{\prime}=\delta_{v,v^{\prime}}v$ for all $v,v^{\prime}\in E^{0}$,
  2. (E1)
    $s(e)e=er(e)=e$ for all $e\in E^{1}$,
  3. (E2)
    $r(e)e^{*}=e^{*}s(e)=e^{*}$ for all $e\in E^{1}$,
  4. (CK1)
    $e^{*}f=\delta_{e,f}r(e)$ for all $e,f\in E^{1}$, and
  5. (CK2)
    $v=\sum_{\{e\in E^{1}|s(e)=v\}}ee^{*}$ for every $v\in{\rm Reg}(E)$.
Let $E$ be a graph. For each $e\in E^{1}$, we call $e^{*}$ the ghost edge. Also, the notation $\mu^{*}=e_{n}^{*}\dots e_{2}^{*}e_{1}^{*}$ stands for the corresponding ghost path of the path $\mu=e_{1}\dots e_{n}$.
2.2. Lie and Jordan structures of associative algebras
Every associative algebra $\mathcal{A}$ over a field $K$ is endowed with a Lie algebra structure for the Lie bracket $[a,b]=ab-ba$ for all $a,b\in\mathcal{A}$. We denote this Lie algebra by $\mathcal{A}^{-}$. For $K$-subspaces $S$ and $T$ of $\mathcal{A}$, we defined $[S,T]={\rm span}_{K}\{ab-ba:a\in S,b\in T\}$. A $K$-subspace $\mathcal{L}$ of $\mathcal{A}$ is called a Lie subalgebra of $\mathcal{A}^{-}$ if $[a,b]\in\mathcal{L}$ for all $a,b\in\mathcal{L}$. For the Lie subalgebra $\mathcal{L}$ of $\mathcal{A}^{-}$, we define inductively the two series
$\mathcal{L}^{(0)}=\mathcal{L},\;\mathcal{L}^{(1)}=[\mathcal{L},\mathcal{L}],\;\dots,\;\mathcal{L}^{(n)}=[\mathcal{L}^{(n-1)},\mathcal{L}^{(n-1)}],\dots$
and
$\mathcal{L}^{0}=\mathcal{L},\;\mathcal{L}^{0}=[\mathcal{L},\mathcal{L}],\;\dots,\;\mathcal{L}^{n}=[\mathcal{L}^{n-1},L],\dots$
The Lie subalgebra $\mathcal{L}$ is called Lie solvable (resp., Lie nilpotent) if there exists $n\geq 0$ (resp., $m\geq 0$) such that $\mathcal{L}^{(n)}=0$ (resp., $\mathcal{L}^{m}=0$). If such the smallest $n$ (resp., $m$) exists, we will say that $\mathcal{L}$ is Lie solvable of index $n$ (resp., Lie nilpotent of index $m$).
Similarly, a Jordan product can be defined on $\mathcal{A}$ by $a\circ b=ab+ba$, for all $a,b\in\mathcal{A}$. If $\mathcal{A}$ is viewed as a Jordan algebra for the product $\circ$, we denote it by $\mathcal{A}^{\circ}$. Note that in the case ${\rm char}(K)=2$, the two structures $\mathcal{A}^{\circ}$ and $\mathcal{A}^{-}$ coincides. A $K$-subspace $\mathcal{J}$ of $\mathcal{A}$ is called a Jordan subalgebra of $\mathcal{A}^{\circ}$ if $a\circ b\in\mathcal{J}$ for all $a,b\in\mathcal{J}$. For two $K$-subspaces $S$ and $T$ of $\mathcal{A}$, we defined
$S\circ T={\rm span}_{K}\{ab+ba:a\in S,b\in T\}.$
For the Jordan subalgebra $\mathcal{J}$ of $\mathcal{A}^{\circ}$, we define the analogous series for $\mathcal{J}$ as follows:
$\mathcal{J}_{(0)}=\mathcal{J},\;\mathcal{J}_{(1)}=\mathcal{J}\circ\mathcal{J},\;\dots,\;\mathcal{J}_{(n)}=\mathcal{J}_{(n-1)}\circ\mathcal{J}_{(n-1)},\dots$
and
$\mathcal{J}_{0}=\mathcal{J},\;\mathcal{J}_{1}=\mathcal{J}\circ\mathcal{J},\;\dots,\;\mathcal{J}_{n}=\mathcal{J}_{n-1}\circ\mathcal{J},\dots$
The Jordan subalgebra $\mathcal{J}$ is called Jordan solvable (resp., Jordan nilpotent) if there exists $n\geq 0$ (resp., $m\geq 0$) such that $\mathcal{J}_{(n)}=0$ (resp., $\mathcal{J}_{m}=0$). The index of a solvable Jordan algebra is defined similarly.
In this paper, we particularly consider the special case where $\mathcal{A}=L_{K}(E)$, $\mathcal{L}=\mathbf{K}_{L_{K}(E)}$ and $\mathcal{J}=\mathbf{S}_{L_{K}(E)}$ with respect to the standard involution on $L_{K}(E)$.
2.3. Homomorphisms of algebras with involutions
Let $K$ be a field, and $(\mathcal{A},^{\star})$ and $(\mathcal{B},^{\natural})$ be $K$-algebras with involutions. A homomorphism of $K$-algebras with involution $\varphi:(\mathcal{A},^{\star})\to(\mathcal{B},^{\natural})$ is a $K$-algebra homomorphism $\varphi:\mathcal{A}\to\mathcal{B}$ such that $(\varphi(ab))^{\natural}=\varphi(b^{\star})\varphi(a^{\star})$.
It is obvious that if $(\mathcal{A},^{\star})$ is isomorphic to $(\mathcal{B},^{\natural})$ as $K$-algebras with involution, then the Lie subalgebras $\mathbf{K}_{\mathcal{A}}$ and $\mathbf{K}_{\mathcal{B}}$, with the respect involutions and $\natural$ respectively, are also isomorphic as Lie algebras over $K$. Similarly, the Jordan subalgebras $\mathbf{S}_{\mathcal{A}}$ and $\mathbf{S}_{\mathcal{B}}$ of $\mathcal{A}^{\circ}$ and $\mathcal{B}^{\circ}$, with the respect involutions and $\natural$ respectively, are also isomorphic as Jordan algebras over $K$.
3. Lie Solvability and Jordan Solvability of $\mathbb{M}_{n}(\mathcal{A})$
Let $K$ be a field, and $\mathcal{A}$ a unital $K$-algebra with involution . Let $\mathbb{M}_{n}(\mathcal{A})$ be the matrix ring of degree $n\geq 1$ over $\mathcal{A}$. Then, it is easy to check that is extended to an involution, which is again denoted by , on $\mathbb{M}_{n}(\mathcal{A})$ by letting
(3.1) $(a_{ij})_{n\times n}^{\natural}=(a_{ji}^{\natural})_{n\times n}$
for every matrix $(a_{ij})\in\mathbb{M}_{n}(\mathcal{A})$. Let $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ be the Lie subalgebra of $\mathbb{M}_{n}(\mathcal{A})^{-}$ with respect to the involution $\natural$ on $\mathbb{M}_{n}(\mathcal{A})$ as defined in (3.1). In this section, we gives conditions on $n$ and $\mathcal{A}$ for $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ to be Lie solvable. First, we fix the following notation: In $\mathbb{M}_{n}(\mathcal{A})$, we denote by $E_{ij}$ the matrix whose $(i,j)$-entry is $1$ and the other entries are 00. The following proposition is the main result of this section:
Proposition 3.1.
Let $K$ be a field, and $\mathcal{A}$ be a unital $K$-algebra with an involution . Assume further that $\mathcal{A}$ is an integral domain. Let $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ be as defined above. Then, the following assertions holds:
  1. (a)
    $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ is not Lie solvable when $n\geq 3$.
  2. (b)
    If ${\rm char}(K)=2$, then $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is not Lie nilpotent.
  3. (c)
    If ${\rm char}(K)=2$, then $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is Lie solvable, with index at most $3$. If moreover, the involution on $\mathcal{A}$ is trivial, then the index is $2$.
  4. (d)
    Assume that ${\rm char}(K)\neq 2$. If the involution on $\mathcal{A}$ is trivial, then $[\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})},\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}]=0$; while if the involution on $\mathcal{A}$ is non-trivial, then $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is not Lie solvable.
Proof.
(a): Assume first that $n\geq 3$. Let $a,b,c\in K$ be such that $c\neq 0$ and $a^{2}+b^{2}\neq 0$. We construct inductively a family of matrices $\{A_{m},B_{m},X_{m}\;|\;m=1,2,\dots\}$ in $\mathbb{M}_{n}(\mathcal{A})$ as follows: Put
$A_{1}=a(E_{12}-E_{21})+b(E_{13}-E_{31}),~{}~{}B_{1}=c(E_{23}-E_{32}),~{}~{}X_{1}=[A_{1},B_{1}];$
and for each $m\geq 2$, we put
$A_{m}=[X_{m-1},B_{m-1}],~{}~{}B_{m}=[X_{m-1},A_{m-1}],\text{ and }X_{m}=[A_{m},B_{m}].$
Since $A_{1},B_{1}\in\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$, we get that $X_{1}\in\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}^{(1)}$. Inductively, it can be checked that $A_{m},B_{m}\in\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}^{(m-1)}$, from which it follows that $X_{m}\in\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}^{(m)}$. For each $m\geq 1$, we claim that there exist $a_{m},b_{m},c_{m}\in K$ such that $c_{m}\neq 0$, $a_{m}^{2}+b_{m}^{2}\neq 0$, and
$A_{m}=a_{m}(E_{12}-E_{21})+b_{m}(E_{13}-E_{31}),$
$B_{m}=c_{m}(E_{23}-E_{32}),$
$X_{m}\neq 0.$
The argument proceeds by induction on $m$. For $m=1$, then we set $a_{1}=a$, $b_{1}=b$ and $c_{1}=c$. Then, we have
$A_{1}=a_{1}(E_{12}-E_{21})+b_{1}(E_{13}-E_{31}),$
$B_{1}=c_{1}(E_{23}-E_{32}),$
$X_{1}=-b_{1}c_{1}(E_{12}-E_{21})+a_{1}c_{1}(E_{13}-E_{31})\neq 0.$
Thus, the claim is true for $m=1$. Now, assume that the claim is true for all $m\geq i$ for some $i\geq 2$. We will prove the claim is also true for $m=i+1$. Now, the inductive assumption shows that there exist $a_{i},b_{i},c_{i}\in K$ such that $c_{i}\neq 0,a_{i}^{2}+b_{i}^{2}\neq 0$, and
$A_{i}=a_{i}(E_{12}-E_{21})+b_{i}(E_{13}-E_{31}),$
$B_{i}=c_{i}(E_{23}-E_{32}),$
$X_{i}=[A_{i},B_{i}]=-b_{i}c_{i}(E_{12}-E_{21})+a_{i}c_{i}(E_{13}-E_{31})\neq 0.$
From this, a direct calculation shows that
$A_{i+1}=[X_{i},B_{i}]=-a_{i}c_{i}^{2}(E_{12}-E_{21})-b_{i}c_{i}^{2}(E_{13}-E_{31}),$
$B_{i+1}=[X_{i},A_{i}]=(a_{i}^{2}+b_{i}^{2})c_{i}(E_{23}-E_{32}),$
$X_{i+1}=[A_{i+1},B_{i+1}]=b_{i}(a_{i}^{2}+b_{i}^{2})c_{i}^{3}(E_{12}-E_{21})-a_{i}(a_{i}^{2}+b_{i}^{2})c_{i}^{3}(E_{13}-E_{31}).$
Since $a_{i}^{2}+b_{i}^{2}\neq 0$, we have $X_{i+1}\neq 0$. The claim is proved. It follows that $X_{m}\neq 0$ for all $m\geq 1$, and so $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}^{(m)}\neq 0$ for all $m\geq 1$. Therefore $\mathbf{K}_{\mathbb{M}_{n}(K)}$ is not Lie solvable, proving (a).
To prove the remaining assertions, we assume from now on that $n=2$. Let us calculate $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ more explicitly. It is obvious that
$\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}=\left\{\begin{pmatrix}a&b\\ -b^{\natural}&c\end{pmatrix}\;|\;a,b,c\in\mathcal{A}\text{ and }a=-a^{\natural},c=-c^{\natural}\right\}.$
Let $A_{i},B_{i}$ ($1\leq i\leq 4$) be arbitrary matrices in $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$. Then, these matrices can be written as
$A_{i}=\left(\begin{array}[]{cc}a_{i}&b_{i}\\ -b_{i}^{\natural}&c_{i}\end{array}\right)\text{ and }B_{i}=\left(\begin{array}[]{cc}u_{i}&v_{i}\\ -v_{i}^{\natural}&w_{i}\end{array}\right),$
where $a_{i},b_{i},c_{i},u_{i},v_{i},w_{i}\in\mathcal{A}$ and $a_{i}=-a_{i}^{\natural}$, $c_{i}=-c_{i}^{\natural}$, $u_{i}=-u_{i}^{\natural}$, $w_{i}=-w_{i}^{\natural}$.
For each $1\leq i\leq 4$, put $X_{i}=[A_{i},B_{i}]$. Because $a_{i}=-a_{i}^{\natural}$, $c_{i}=-c_{i}^{\natural}$, $u_{i}=-u_{i}^{\natural}$, and $w_{i}=-w_{i}^{\natural}$, we get
(3.2) $X_{i}=\left(\begin{array}[]{cc}r_{i}&s_{i}\\ -s_{i}^{\natural}&-r_{i}\end{array}\right),\text{ where }$
(3.3) $r_{i}=b_{i}^{\natural}v_{i}-b_{i}v_{i}^{\natural}\text{ and }s_{i}=v_{i}(a_{i}-c_{i})+b_{i}(w_{i}-u_{i}).$
It follows that
(3.4) $[X_{1},X_{2}]=\left(\begin{array}[]{cc}s_{2}s_{1}^{\natural}-s_{1}s_{2}^{\natural}&0\\[8.5359pt] 2(r_{1}s_{2}^{\natural}-s_{1}^{\natural}r)&s_{2}^{\natural}s_{1}-s_{1}^{\natural}s_{2}\end{array}\right),\text{ where }$
(3.5) $s_{2}s_{1}^{\natural}-s_{1}s_{2}^{\natural}\hskip 2.84544pt=(v_{2}v_{1}^{\natural}-v_{1}v_{2}^{\natural})(a_{1}-c_{1})(c_{2}-a_{2})$
$\hskip 54.06006pt+(v_{2}b_{1}^{\natural}-b_{1}v_{2}^{\natural})(a_{2}-c_{2})(u_{1}-w_{1})$
$\hskip 54.06006pt+(v_{1}b_{2}^{\natural}-b_{2}v_{1}^{\natural})(w_{2}-u_{2})(a_{1}-c_{1})$
$\hskip 54.06006pt+(b_{1}b_{2}^{\natural}-b_{2}b_{1}^{\natural})(u_{1}-w_{1})(u_{2}-w_{2}).$
(b) and (c): Assume that ${\rm char}(K)=2$. As $02(r_{1}s_{2}^{\natural}-s_{1}^{\natural}r)=$ 0, it follows from (3.4) that $[X_{1},X_{2}]$ is a diagonal matrix. Similarly, we also obtain that $[X_{3},X_{4}]$ is also a diagonal matrix. Thus $[[X_{1},X_{2}],[X_{3},X_{4}]]=0$, which means that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(3)}=0$. We divide our situation into two possible cases:
Case 1. ${\rm char}(K)=2$ and the involution on $\mathcal{A}$ is trivial. As $b_{i}^{\natural}=b_{i}$ and $v_{i}^{\natural}=v_{i}$, we get from (3.3) that $r_{i}=0$. It follows from (3.2) that $X_{i}=\left(\begin{array}[]{cc}0&s_{i}\\ -s_{i}^{\natural}&0\end{array}\right)$, and so $[X_{1},X_{2}]=0$; or equivalently, we have $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(2)}=0$. In other words, $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is Lie solvable of index 2. This completes the proof of the second assertion of (c).
Case 2. ${\rm char}(K)=2$ and the involution on $\mathcal{A}$ is non-trivial. In this case, suitable choices of the matrices $A_{1},A_{2},B_{1},B_{2}$ yield that $[X_{1},X_{2}]\neq 0$. For example, if we choose $a_{1}=c_{1}$, $a_{2}=c_{2}$, $b_{1}=1$, $u_{1}\neq w_{1}$, $w_{1}=w_{2}=0$, $0\neq u_{1}=u_{2}\in K$, $b_{2}\neq b_{2}^{\natural}$, then it follows from (3.5) that $s_{2}s_{1}^{\natural}-s_{1}s_{2}^{\natural}=(b_{2}^{\natural}-b_{2})u_{1}^{2}\neq 0$. Therefore, we get from (3.4) that $[X_{1},X_{2}]\neq 0$. It follows that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is Lie solvable of index 3. This completes the proof of (c).
Now, we give the proof of (b). Let $A=E_{12}+E_{21}$ and $B=E_{11}$, which belong to $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$. Define
[…]
(d): Recall that
$\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}=\left\{\begin{pmatrix}a&b\\ -b^{\natural}&c\end{pmatrix}\;|\;a,b,c\in\mathcal{A}\text{ and }a=-a^{\natural},c=-c^{\natural}\right\}.$
We consider the following two possible cases:
Case 3. ${\rm char}(K)\neq 2$ and the involution on $\mathcal{A}$ is trivial. Because $a^{\natural}=a$, the equation $a=-a^{\natural}$ is equivalent to $02a=$ 0, and hence $a=0$. Similarly, we also obtain that $c=0$. Moreover, as $b^{\natural}=b$, we finally get that
$\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}=\left\{\begin{pmatrix}0&b\\ -b&0\end{pmatrix}\;|\;b\in\mathcal{A}\right\}.$
It follows that $[\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})},\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}]=0$.
Case 4. ${\rm char}(K)\neq 2$ and the involution on $\mathcal{A}$ is non-trivial. Fix $a\in\mathcal{A}$ such that $a\neq a^{\natural}$. Put $u=a-a^{\natural}$. Then, we have $u=-u^{\natural}$. We construct inductively a family of matrices $A_{m}$, $B_{m}$, and $X_{m}$ as follows. For $m=1$, we put
$A_{1}=\left(\begin{array}[]{cc}0&u\\ u&0\end{array}\right),~{}B_{1}=\left(\begin{array}[]{cc}u&0\\ 0&-u\end{array}\right)\text{ and }X_{1}=[A_{1},B_{1}],$
For $m\geq 2$, we put
$A_{m}=[X_{m-1},B_{m-1}],B_{m}=[X_{m-1},A_{m-1}]\text{ and }X_{m}=[A_{m},B_{m}].$
Since $A_{1},B_{1}\in\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$, we get that $X_{1}\in\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(1)}$. Inductively, it can be checked that $A_{m},B_{m}\in\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(m-1)}$, from which it follows that $X_{m}\in\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(m)}$. We claim that for each positive integer $m$, there exists $v_{m}\in\mathcal{A}$ such that $0\neq v_{m}=-v_{m}^{\natural}$ and
$A_{m}=\left(\begin{array}[]{cc}0&v_{m}\\ v_{m}&0\end{array}\right),$
$B_{m}=\left(\begin{array}[]{cc}(-1)^{m+1}v_{m}&0\\ 0&(-1)^{m}v_{m}\end{array}\right),$
$X_{m}=\left(\begin{array}[]{cc}0&(-1)^{m}2v_{m}^{2}\\ (-1)^{m+1}2v_{m}^{2}&0\end{array}\right)\neq 0.$
We proceed by induction on $m$. For $m=1$, we set $v_{1}=u$. Then, we have $0\neq v_{1}=v_{1}^{\natural}$ and $X_{1}=[A_{1},B_{1}]=\left(\begin{array}[]{cc}0&-2u^{2}\\ 2u^{2}&0\end{array}\right)\neq 0$. Thus, the claim is true for the case $m=1$. Now, assume that the claim is true for $m=i\geq 2$. That means there exist $v_{i}\in\mathcal{A}$ such that $0\neq v_{i}=-v_{i}^{\natural}$ and
$A_{i}=\left(\begin{array}[]{cc}0&v_{i}\\ v_{i}&0\end{array}\right),$
$B_{i}=\left(\begin{array}[]{cc}(-1)^{i+1}v_{i}&0\\ 0&(-1)^{i}v_{i}\end{array}\right),$
$X_{i}=\left(\begin{array}[]{cc}0&(-1)^{i}2v_{i}^{2}\\ (-1)^{i+1}2v_{i}^{2}&0\end{array}\right)\neq 0.$
Now, we prove the claim for the case $m=i+1$. A straightforward calculation shows that
$A_{i+1}=\left(\begin{array}[]{cc}0&4v_{i}^{3}\\ 4v_{i}^{3}&0\end{array}\right),$
$B_{i+1}=\left(\begin{array}[]{cc}(-1)^{i}4v_{i}^{3}&0\\ 0&(-1)^{i+1}4v_{i}^{3}\end{array}\right),$
$X_{i+1}=\left(\begin{array}[]{cc}0&(-1)^{i+1}32v_{i}^{6}\\ (-1)^{i+2}32v_{i}^{6}&0\end{array}\right)\neq 0.$
Now, put $v_{i+1}=4v_{i}^{3}$. Then, we have
$v_{i+1}^{\natural}=(4v_{i}^{3})^{\natural}=4(v_{i}^{\natural})^{3}=-4v_{i}^{3}=-v_{i+1}\neq 0.$
Therefore, the claim is true for $m=i+1$. The claim is proved. It follows that $X_{m}\neq 0$ for all $m\geq 1$. In other words, we have $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(m)}$ for all $m\geq 1$, which means that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is not Lie solvable. ∎
In the remainder of this section, we record two corollaries of Proposition 3.1 for further use. In the proposition, let $A=K$ and be the trivial involution on $K$. Then, the extended involution on $\mathbb{M}_{n}(K)$ given by (3.1) is just the transpose operation on $\mathbb{M}_{n}(K)$; that is, for every matrix $(a_{ij})_{n\times n}\in\mathbb{M}_{n}(K)$, we have
(3.6) $\left((a_{ij})_{n\times n}\right)^{\natural}=\left(a_{ji}\right)_{n\times n}.$
With respect to this involution, the Lie algebra $\mathbf{K}_{\mathbb{M}_{n}(K)}$ consists precisely of skew-symmetric matrices in $\mathbb{M}_{n}(K)$:
$\mathbf{K}_{\mathbb{M}_{n}(K)}=\left\{(a_{ij}):a_{ji}=-a_{ij}\text{ for all }1\leq i,j\leq n\right\}.$
As a consequence of Proposition 3.1, we easily obtain the following:
Corollary 3.2.
Let $K$ be a field, and $n\geq 2$ an integer. Let $\mathbf{K}_{\mathbb{M}_{n}(K)}$ be the Lie subalgebra of skew-symmetric matrices of $\mathbb{M}_{n}(K)^{-}$. The following assertions hold:
  1. (1)
    Suppose that $n=2$. If ${\rm char}(K)\neq 2$, then $[\mathbf{K}_{\mathbb{M}_{n}(K)},\mathbf{K}_{\mathbb{M}_{n}(K)}]=0$; and, if ${\rm char}(K)=2$, then $\mathbf{K}_{\mathbb{M}_{n}(K)}$ is Lie solvable of index $2$ but not Lie nilpotent.
  2. (2)
    If $n\geq 3$, then $\mathbf{K}_{\mathbb{M}_{n}(K)}$ is not Lie solvable.
For any field $K$, let $K[x,x^{-1}]$ be the Laurent polynomial $K$-algebra in $x$; that is,
$K[x,x^{-1}]=\left\{\sum_{i=k}^{n}a_{i}x^{i}\;|\;a_{i}\in K\text{ and }k\geq n\in\mathbb{Z}\right\}.$
Then, the map ${}^{\natural}:K[x,x^{-1}]\to K[x,x^{-1}]$ defined by $(\sum_{i}a_{i}x^{i})^{\natural}=\sum_{i}a_{i}x^{-i}$ is a non-trivial involution on $K[x,x^{-1}]$. Now, in Proposition 3.1, we consider the case where $\mathcal{A}=K[x,x^{-1}]$ and extend the involution on $K[x,x^{-1}]$ to the matrix ring $\mathbb{M}_{n}(K[x,x^{-1}])$ as defined in (3.1). This extended involution ${}^{\natural}:\mathbb{M}_{n}(K[x,x^{-1}])\to\mathbb{M}_{n}(K[x,x^{-1}])$ is precisely given by the following rule:
(3.7) $\left((f_{ij}(x))_{n\times n}\right)^{\natural}=\left(f_{ji}(x^{-1})\right)_{n\times n}$
for each matrix $(f_{ij}(x))\in\mathbb{M}_{n}(K[x,x^{-1}])$. With respect to this involution, we easily get the following corollary to Proposition 3.1:
Corollary 3.3.
Let $K[x,x^{-1}]$ be the Laurent polynomial $K$-algebra. Let $\mathbf{K}_{\mathbb{M}_{n}(K[x,x^{-1}])}$ be the Lie subalgebra of skew-symmetric elements of $\mathbb{M}_{n}(K[x,x^{-1}])^{-}$ with respect to the involution defined by ((((3.7)))). Then, the following assertions hold:
  1. (1)
    If $\mathrm{char}(K)=2$, then $\mathbf{K}_{\mathbb{M}_{2}(K[x,x^{-1}])}$ is Lie solvable of index $3$ but not Lie nilpotent.
  2. (2)
    If $\mathrm{char}(K)\neq 2$, then $\mathbf{K}_{\mathbb{M}_{2}(K[x,x^{-1}])}$ is not Lie solvable.
4. Lie Solvability and Jordan Solvability of $L_{K}(E)$
Throughout of this section, we will denote by $\mathbf{K}_{L_{K}(E)}$ and $\mathbf{S}_{L_{K}(E)}$ the Lie subalgebra of $L_{K}(E)^{-}$ and the Jordan subalgebra of $L_{K}(E)^{\circ}$with respect to standard involution on $L_{K}(E)$ as defined before, respectively. The main aim of this section is to provide a necessary and sufficient condition on the graph $E$ and the field $K$ for $\mathbf{K}_{L_{K}(E)}$ and $\mathbf{S}_{L_{K}(E)}$ to be Lie and Jordan solvable, respectively. We start this section a lemma which determines a $K$-basis of $L_{K}(E)$:
Lemma 4.1 ([1, Corollary 1.5.12]).
Let $E$ be a graph and $K$ a field. Put $\mathscr{A}=\{\lambda\nu^{*}\;|\;\lambda,\nu\in{\rm Path}(E)\text{ and }r(\lambda)=r(\nu)\}$. For each $v\in{\rm Reg}(E)$, let $\{e_{1}^{v},\dots,e_{n_{v}}^{v}\}$ be an enumeration of the elements of $s^{-1}(v)$. Then, a $K$-basic of $L_{K}(E)$ is given by the family
$\mathscr{B}=\mathscr{A}\backslash\{\lambda e_{n_{v}}^{v}(e_{n_{v}}^{v})^{*}\nu^{*}\;:\;r(\lambda)=r(\nu)=v\in{\rm Reg}(E)\}.$
The next lemma provides a necessary condition for $\mathbf{K}_{L_{K}(E)}$ to be Lie solvable. The ideas used to prove of this lemma come from the proof of [12, Lemma 2.2 and Theorem 2.3]. It can be said that, in $L_{K}(E)$, we can easily choose a suitable subalgebra which is isomorphic to the matrix ring $\mathbb{M}_{3}(K)$. Then, Corollary 3.2 is applied to get all possible cases for $\mathbf{K}_{L_{K}(E)}$ to be Lie solvable. For convenience of the readers, we still include the proof here.
Lemma 4.2.
. Let $L_{K}(E)$ be the Leavitt path algebra of a graph $E$ with coefficients in a field $K$. Then, $\mathbf{K}_{L_{K}(E)}$ is not Lie solvable if one of the following cases occurs:
  1. (1)
    $E$ contains a cycle with an exit.
  2. (2)
    $E$ contains one of following subgraphs:
[…]
Proof.
(1) Assume that $E$ contains a cycle $c$ with an exit $f$. Then, according to Lemma 4.1, the set $\{c^{m}ff^{*}(c^{*})^{n}:1\leq m,n\leq 3\}$ is linearly independent over $K$ satisfying the rule
$(c^{m}ff^{*}(c^{*})^{i})(c^{j}ff^{*}(c^{*})^{n})=\delta_{ij}c^{m}ff^{*}(c^{*})^{n},$
where $\delta_{ij}$ is the Kronecker delta. It follows that the $K$-subalgebra of $L_{K}(E)$, say $A$, generated by $\{c^{m}ff^{*}(c^{*})^{n}:1\leq m,n\leq 3\}$ is isomorphic to $\mathbb{M}_{3}(K)$. The restriction of the standard involution on $L_{K}(E)$ to $A$ is still an involution on $A$ which satisfies $\mathbf{K}_{A}\subseteq\mathbf{K}_{L_{K}(E)}$. Let be the involution on $\mathbb{M}_{3}(K)$ as defined in (3.6). Then, it can be checked that ($A,^{\star}$) is isomorphic to $(\mathbb{M}_{3}(K),^{\natural})$ as $K$-algebras with involution. Thus, we have that $\mathbf{K}_{A}$ is isomorphic to $\mathbf{K}_{\mathbb{M}_{3}(K)}$ as Lie algebras over $K$. Because $\mathbf{K}_{\mathbb{M}_{3}(K)}$ is not Lie solvable (see Corollary 3.2), it follows that $\mathbf{K}_{A}$, and hence $\mathbf{K}_{L_{K}(E)}$ is not Lie solvable.
(2) For the proof of this assertion, we assume first that $E$ contains $F_{1}$. Then, $v,f,ef$ are all paths ending at $v$. It is straightforward to check the subalgebra $B$ of $L_{K}(E)$ generated by $v$, $f$, $ef$, $f^{*}$, $f^{*}e^{*}$, $ff^{*}$, $eff^{*}$, $ff^{*}e^{*}$, $eff^{*}e^{*}$ is isomorphic to $\mathbb{M}_{3}(K)$ via map: $eff^{*}e^{*}\mapsto E_{11}$, $ff^{*}\mapsto E_{22}$, $v\mapsto E_{33}$, $eff^{*}\mapsto E_{12}$, $ff^{*}e^{*}\mapsto E_{21}$, $f\mapsto E_{23}$, $f^{*}\mapsto E_{32}$, $ef\mapsto E_{13}$, and $f^{*}e^{*}\mapsto E_{31}$ (see also a similar techniques used in the proof of [1, Lemma 2.6.4]). The restriction of the involution to $B$ is also an involution on $B$ for which $\mathbf{K}_{B}\subseteq\mathbf{K}_{L_{K}(E)}$. Thus, $\mathbf{K}_{B}$ is isomorphic to $\mathbf{K}_{\mathbb{M}_{3}(K)}$ as Lie algebras over $K$. Therefore, by Corollary 3.2, we conclude that $\mathbf{K}_{B}$, and so $\mathbf{K}_{L_{K}(E)}$ is not Lie solvable.
Finally, assume that $E$ contains $F_{2}$ or $F_{3}$. Then, the subalgebra $C$ generated by $v,e,f,e^{*},f^{*},ee^{*},ff^{*},ef^{*},fe^{*}$ is isomorphic to $\mathbb{M}_{3}(K)$ via map $ee^{*}\mapsto E_{11}$, $ff^{*}\mapsto E_{22}$, $v\mapsto E_{33}$, $ef^{*}\mapsto E_{12}$, $fe^{*}\mapsto E_{21}$, $e\mapsto E_{13}$, $e^{*}\mapsto E_{31}$, $f\mapsto E_{23}$, $f^{*}\mapsto E_{32}$. A similar argument as in previous paragraph shows that $\mathbf{K}_{L_{K}(E)}$ is not Lie solvable too. The proof is now completed. ∎
As a consequence of Lemma 4.2, we have the following remark.
Remark 1.
Let $L_{K}(E)$ be the Leavitt path algebra of a graph $E$ with coefficients in a field $K$. If $\mathbf{K}_{L_{K}(E)}$ is Lie solvable then $E$ is a disjoint union of following graphs:
$E_{1}:$$\bullet$
$E_{2}:$$\bullet$
[…]
in which $\{u_{i}|i\in I\}$ and $\{v_{j}|j\in J\}$, where $I$ and $J$ are indexing non-empty sets of arbitrary cardinality.
Remark 2.
Let $E_{4},E_{5},E_{6}$ be graphs as defined above. If $m:=|I|$ and $n:=|J|$ are finite, then by [1, Corollary 2.7.5], we get that
$L_{K}(E_{1})\cong K,\;\;L_{K}(E_{2})\cong K[x,x^{-1}],\;\;L_{K}(E_{3})\cong\mathbb{M}_{2}(K[x,x^{-1}]),$
$L_{K}(E_{4})\cong\bigoplus_{i=1}^{n}\mathbb{M}_{2}(K),\;\;L_{K}(E_{5})\cong\bigoplus_{i=1}^{m}\mathbb{M}_{2}(K[x,x^{-1}]),$
$L_{K}(E_{6})\cong\bigoplus_{i=1}^{n}\mathbb{M}_{2}(K)\oplus\bigoplus_{i=1}^{m}\mathbb{M}_{2}(K[x,x^{-1}]).$
In what follows, we give a sufficient condition for $\mathbf{K}_{L_{K}(E)}$ to be Lie solvable. We proceed by proving several auxiliary lemmas.
Lemma 4.3.
Let $\mathcal{A}$ be a $K$-algebra with an involution , and $\mathbf{J}$ be an indexing set of arbitrary cardinality. Let $I_{j},j\in\mathbf{J}$, be ideals of $\mathcal{A}$ satisfying the property that for each $j$, if $a\in I_{j}$ then $a^{\star}\in I_{j}$. Put $M=\sum_{j\in\mathbf{J}}I_{j}$. If $M=\bigoplus_{j\in\mathbf{J}}I_{j}$, then $\mathbf{K}_{M}=\bigoplus_{j\in\mathbf{J}}\mathbf{K}_{I_{j}}$. Consequently, $\mathbf{K}_{M}$ is Lie solvable ((((nilpotent)))) if and only if $\mathbf{K}_{I_{j}}$ is Lie solvable ((((nilpotent)))) of a common index for all $j\in\mathbf{J}$.
Proof.
Let $x\in\mathbf{K}_{M}$. Then, we may write $x=a_{j_{1}}+a_{j_{2}}+\dots+a_{j_{n}}$, where $a_{j_{k}}\in I_{j_{k}}$ for all $1\leq k\leq n$. The equation $x^{\star}=-x$ implies that
$a_{j_{1}}+a_{j_{1}}^{\star}=-(a_{j_{2}}+a_{j_{2}}^{\star}+\cdots+a_{j_{n}}+a_{j_{n}}^{\star}).$
Because $a_{j_{k}}\in I_{j_{k}}$ for all $i$, the above equation shows that $a_{j_{1}}+a_{j_{1}}^{\star}\in I_{j_{1}}\cap(I_{j_{2}}\oplus\cdots\oplus I_{j_{n}})=0$. It follows that $a_{1}+a_{1}^{\star}=0$, and so $a_{j_{1}}\in\mathbf{K}_{I_{j_{1}}}$. Similarly, we also obtain that $a_{i}\in\mathbf{K}_{I_{j_{k}}}$ for $k\in\{2,\dots,n\}$. In other words, we have $\mathbf{K}_{M}\subseteq\sum_{j\in\mathbf{J}}\mathbf{K}_{I_{j}}=\bigoplus_{j\in\mathbf{J}}\mathbf{K}_{I_{j}}$. The inverse inclusion is clear. The final assertion follows from the fact that $\mathbf{K}_{M}^{(n)}=\bigoplus_{j\in\mathbf{J}}\mathbf{K}_{I_{j}}^{(n)}$ and $\mathbf{K}_{M}^{n}=\bigoplus_{j\in\mathbf{J}}\mathbf{K}_{I_{j}}^{n}$ for all integer $n\geq 1$. ∎
Lemma 4.4.
Let $K$ be a field, and $E_{1},E_{2},E_{3}$ be graphs as defined in Remark 1. Then, the following assertions hold:
  1. (1)
    $\mathbf{K}_{L_{K}(E_{1})}$ and $\mathbf{K}_{L_{K}(E_{2})}$ are Lie solvable of index $1$.
  2. (2)
    $\mathbf{K}_{L_{K}(E_{3})}$ is Lie solvable ((((of index $3$)))) if and only if $\mathrm{char}(K)=2$. Additionally, $\mathbf{K}_{L_{K}(E_{3})}$ is not Lie nilpotent.
Proof.
The assertion (i) follows immediately from the fact that $L_{K}(E_{1})$ and $L_{K}(E_{2})$ are commutative. Moreover, because $L_{K}(E_{3})\cong\mathbb{M}_{2}(K[x,x^{-1}])$, we get that $\mathbf{K}_{L_{K}(E_{3})}$ is not Lie nilpotent. ∎
Lemma 4.5.
Let $K$ be a field, and $E_{4},E_{5},E_{6}$ be graphs as defined in Remark 1. Then, the following assertions hold:
  1. (1)
    If $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable, then it is Lie solvable of index $l\leq 3$. More precisely, if $|I|$ is finite and ${\rm char}(K)\neq 2$ ((((resp., ${\rm char}(K)=2$)))), then $l=1$ ((((resp., $l=2$)))). If $|I|$ is infinite and ${\rm char}(K)\neq 2$ ((((reps., ${\rm char}(K)=2$)))), then $l=2$ ((((resp., $l=3$)))). Additionally, $\mathbf{K}_{L_{K}(E_{4})}$ is Lie nilpotent if and only if $|I|$ is finite and ${\rm char}(K)\neq 2$, if and only if $[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]=1$.
  2. (2)
    $\mathbf{K}_{L_{K}(E_{5})}$ is Lie solvable ((((of index $l\leq 4$)))) if and only if ${\rm char}(K)=2$. More precisely, if $|J|$ is finite ((((resp., infinite)))), then $l=3$ ((((resp., $l=4$)))). Additionally, $\mathbf{K}_{L_{K}(E_{5})}$ is not Lie nilpotent.
  3. (3)
    $\mathbf{K}_{L_{K}(E_{6})}$ is Lie solvable ((((of index $l\leq 4$)))) if and only if ${\rm char}(K)=2$. More precisely, if both $|I|$ and $|J|$ are finite ((((resp., either $|I|$ or $|J|$ is infinite)))), then $l=3$ ((((resp., $l=4$)))). Additionally, $\mathbf{K}_{L_{K}(E_{6})}$ is not Lie nilpotent.
Proof.
(1) Consider the graph $E_{4}$. For each $1\leq i\leq n$, let $I_{i}$ be the ideal of $L_{K}(E_{4})$ generated by $u_{i}$. Then, it is clear that, for each $i$,
$I_{i}={\rm span}_{K}\{u_{i},e_{i},e_{i}^{*},e_{i}e_{i}^{*}\}.$
Moreover, according to Lemma 4.1, the set $\{u_{i},e_{i},e_{i}^{*},e_{i}e_{i}^{*}\}$ is linearly independent over $K$. Thus, it is can be checked that $I_{i}\cong\mathbb{M}_{2}(K)$ via the mapping $u_{i}\mapsto E_{11}$, $e_{i}e_{i}^{*}\mapsto E_{22}$, $e_{i}\mapsto E_{12}$ and $e_{i}^{*}\mapsto E_{21}$. Moreover, we have $\sum_{i=1}^{n}I_{i}=\bigoplus_{i=1}^{n}I_{i}$. Put $M=\bigoplus_{i=1}^{n}I_{i}$. Then, according to Lemma 4.3 and Corollary 3.2, we conclude that $\mathbf{K}_{M}=\bigoplus_{i=1}^{n}\mathbf{K}_{I_{i}}$, and that $\mathbf{K}_{M}$ is Lie solvable. Moreover, in view of Corollary 3.2, if ${\rm char}(K)\neq 2$, then $[\mathbf{K}_{M},\mathbf{K}_{M}]=0$; and, if ${\rm char}(K)=2$ then $\mathbf{K}_{M}$ is Lie solvable of index $2$ but not Lie nilpotent. Now, we consider two possible cases:
Case 1. $n$ is finite. In this case, as $u\in M$, we get that $L_{K}(E_{4})=M$, from which it follows that $\mathbf{K}_{L_{K}(E_{4})}=\mathbf{K}_{M}$. Thus, all conclusions follow immediately in this case.
Case 2. $n$ is infinite. In this case, $u$ is an infinite emitter and $u\not\in M$. This implies that $L_{K}(E_{4})=Ku\oplus M$ as $K$-vector spaces. It follows that
$[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]\subseteq[L_{K}(E_{4}),L_{K}(E_{4})]=[M,M]\subseteq M.$
As $M$ is Lie solvable, we conclude that $[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]$, and hence $\mathbf{K}_{L_{K}(E_{4})}$, is Lie solvable. To determine the solvable index of $\mathbf{K}_{L_{K}(E_{4})}$, we consider the following subcases:
[…]
Subcase 2.2. ${\rm char}(K)=2$. As $[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]\subseteq\mathbf{K}_{M}$ which is Lie solvable of index $2$, it follows that $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable of index at most $3$. Because $v\in\mathbf{K}_{L_{K}(E_{4})}\backslash\mathbf{K}_{M}$, we conclude that $\mathbf{K}_{L_{K}(E_{4})}\neq[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]$. It follows that $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable of index $3$. On the other hand, as $\mathbf{K}_{I_{i}}$ is not Lie nilpotent and $\mathbf{K}_{I_{i}}\subseteq\mathbf{K}_{L_{K}(E_{4})}$ for all $i$, we obtain that $\mathbf{K}_{L_{K}(E_{4})}$ is not Lie nilpotent. The proof of assertion (1) is complete.
(2) Consider the graph $E_{5}$. For each $1\leq i\leq m$, let $J_{i}$ be the ideal of $L_{K}(E_{5})$ generated by the $v_{j}$’s. Then, we have $J_{j}\cong\mathbb{M}_{2}(K[x,x^{-1}])$ for all $j$, and $\sum_{i=1}^{n}J_{j}=\bigoplus_{j=1}^{n}J_{j}$. Therefore, one can repeat the arguments of the proof of assertion(1), together with using Corollary 3.3 instead of Corollary 3.2, to get (2).
(3) Consider the graph $E_{6}$. Let $I_{i}$ and $J_{j}$ be respectively the ideals of $L_{K}(E_{6})$ generated by $v_{i}$ and $u_{j}$ for $1\leq i\leq n$ and $1\leq j\leq m$. Then, for each $i$ and each $j$, we have $I_{i}\cong\mathbb{M}_{2}(K)$ and $J_{j}\cong\mathbb{M}_{2}(K[x,x^{-1}])$. It can be checked that
$\sum_{i=1}^{n}I_{i}+\sum_{j=1}^{m}J_{j}=\bigoplus_{i=1}^{n}I_{i}\oplus\bigoplus_{j=1}^{m}J_{j}.$
Put $M=\bigoplus_{i=1}^{n}I_{i}$ and $N=\bigoplus_{j=1}^{m}J_{j}$. Then, as before, $\mathbf{K}_{M}$ is Lie solvable of index at most $3$; and, $\mathbf{K}_{N}$ is Lie solvable (of index at most $4$) if and only if ${\rm char}(K)=2$. Moreover, it is clear that $L_{K}(E_{6})=Kw+(M\oplus N)$. As $N$ is not Lie nilpotent, we get that $\mathbf{K}_{L_{K}(E_{6})}$ is not Lie nilpotent. Moreover, we have that $\mathbf{K}_{L_{K}(E_{6})}$ is Lie solvable if and only if both $\mathbf{K}_{M}$ and $\mathbf{K}_{N}$ are Lie solvable, if and only if ${\rm char}(K)=2$. Now, assume that ${\rm char}(K)=2$. To determine the Lie solvability index of $\mathbf{K}_{L_{K}(E_{6})}$, we consider the following cases. If both $|I|$ and $|J|$ are finite, then $w\in M\oplus N$; and so, $L_{K}(E_{6})=M\oplus N$. Again, by Lemma 4.3, we obtain that $\mathbf{K}_{L_{K}(E_{6})}=\mathbf{K}_{M}\oplus\mathbf{K}_{N}$. According to Corollary 3.2, $\mathbf{K}_{M}$ is Lie solvable of index $2$, while by Corollary 3.3, $\mathbf{K}_{N}$ is Lie solvable of index $3$. It follows that $\mathbf{K}_{L_{K}(E_{6})}$ is Lie solvable of index $3$. If either $|I|$ or $|J|$ is infinite, then $w\notin M\oplus N$. Thus, a similar argument as used in Case 2 above shows that $\mathbf{K}_{L_{K}(E_{6})}$ is Lie solvable of index $4$. ∎
We now present the main result of this section: a sufficient and necessary condition for $\mathbf{K}_{L_{K}(E)}$ to be Lie solvable.
Theorem 4.6.
Let $K$ be a field and $E$ a graph. Then, the followings hold:
  1. (1)
    If ${\rm char}(K)=2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((((of index $\leq 4$)))) if and only if $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$. Specially, the index is $4$ precisely when $E$ contains either $E_{5}$ or $E_{6}$ with infinite $|I|$ or $|J|$.
  2. (2)
    If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((((of index $\leq 2$)))) if and only if $E$ is a disjoint union of graphs $E_{1}$, $E_{2}$ and $E_{4}$. Specially, the index is $2$ precisely when $E$ contains $E_{4}$ with infinite $|I|$.
Proof.
To prove (1), we assume first that ${\rm char}(K)=2$. For a proof of the “only if” part, we assume that $L_{K}(E)$ is Lie solvable. Then by Remark 1, the graph $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$. Conversely, we assume that $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$. Then, $L_{K}(E)$ is a direct summand of ideals each of which is isomorphic to $L_{K}(E_{1})$, $\dots$, $L_{K}(E_{6})$. In view of Lemmas 4.4 and 4.5, we get that $L_{K}(E_{1})$, $\dots$, $L_{K}(E_{6})$ are Lie solvable, and so the assertion (1) follows. Finally, the proof of (2) is similar to that of (1). ∎
If $E$ is row-finite, then by Remark 2, we can obtain more detailed descriptions as follows:
Corollary 4.7.
Let $K$ be a field and $E$ a row-finite graph.
  1. (1)
    If ${\rm char}(K)=2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((((of index $\leq 3$)))) if and only if $E$ is a disjoint union graphs $E_{1},\dots,E_{6}$ with finite $|I|$ and $|J|$. In this case,
    $L_{K}(E)\cong\bigoplus_{I_{1}}K\oplus\bigoplus_{I_{2}}K[x,x^{-1}]\oplus\bigoplus_{I_{3}}\mathbb{M}_{2}(K)\oplus\bigoplus_{I_{4}}\mathbb{M}_{2}(K[x,x^{-1}]),$
    for suitable ((((possibly infinite)))) indexing sets $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$. Moreover, the index is $3$ precisely when $E$ contains $E_{3}$, $E_{5}$ or $E_{6}$.
  2. (2)
    If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((((of index $\leq 2$)))) if and only if $E$ is a disjoint union of graphs $E_{1}$, $E_{2}$ and $E_{4}$ with finite $|I|$. In this case, we have
    $L_{K}(E)\cong\bigoplus_{J_{1}}K\oplus\bigoplus_{J_{2}}K[x,x^{-1}]\oplus\bigoplus_{J_{3}}\mathbb{M}_{2}(K),$
    for suitable ((((possibly infinite)))) indexing sets $J_{1}$, $J_{2}$, $J_{3}$. Moreover, the index is $2$ precisely when $E$ contains $E_{3}$.
The following theorem provides a necessary and sufficient condition for $\mathbf{K}_{L_{K}(E)}$ to be Lie nilpotent. The proof of this theorem is similar to that of Theorem 4.6, so it should be omitted.
Theorem 4.8.
Let $K$ be a field and $E$ a graph.
  1. (1)
    If ${\rm char}(K)=2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie nilpotent if and only if $E$ is a disjoint union graphs $E_{1}$ and $E_{2}$.
  2. (2)
    If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie nilpotent if and only if $E$ is a disjoint union graphs $E_{1}$, $E_{2}$, and $E_{4}$ with finite $n$.
In particular, $\mathbf{K}_{L_{K}(E)}$ is Lie nilpotent if and only if $[\mathbf{K}_{L_{K}(E)},\mathbf{K}_{L_{K}(E)}]=0$.
Corollary 4.9.
Let $K$ be a field and $E$ a graph. If $E$ contains an infinite emitter, then $\mathbf{K}_{L_{K}(E)}$ is not Lie nilpotent.
The following corollary addresses a situation where being Lie solvable and being Lie nilpotent are equivalent.
Corollary 4.10.
Let $K$ be a field and $E$ a row-finite graph. If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable if and only if it is Lie nilpotent, if and only if $E$ is a disjoint union of graphs $E_{1}$, $E_{2}$ and $E_{4}$.
Proof.
Assume that $L_{K}(E)$ is Lie solvable. In view of Corollary 4.7, $E$ is a disjoint union of a finite number of graphs $E_{1}$, $E_{2}$ or $E_{4}$ with finite $|I|$. Therefore, $L_{K}(E)$ is Lie nilpotent by Theorem 4.8. ∎
In the remainder of this section, we investigate the solvability of the Jordan algebra $\mathbf{S}_{L_{K}(E)}$.
Theorem 4.11.
Let $K$ be a field and $E$ a graph. Then, the following assertions hold:
  1. (1)
    If ${\rm char}(K)=2$, then $\mathbf{S}_{L_{K}(E)}$ is Jordan solvable ((((resp., nilpotent)))) if and only if $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$ ((((resp., $E_{1}$ and $E_{2}$)))).
  2. (2)
    If ${\rm char}(K)\neq 2$, then $\mathbf{S}_{L_{K}(E)}$ is not Jordan nilpotent.
Proof.
(1) Assume that $\mathrm{char}(K)=2$. Then,
$a\circ b=ab+ba=ab-ba=[a,b]\text{ for any }a,b\in L_{K}(E).$
It follows that $\mathbf{S}_{L_{K}(E)}=\mathbf{K}_{L_{K}(E)}$. Thus, $\mathbf{S}_{L_{K}(E)}$ is Jordan solvable (resp., nilpotent) if and only if $\mathbf{K}_{L_{K}(E)}$ is Lie solvable (resp., nilpotent). By Theorems 4.6 and 4.8, all conclusions of (1) are asserted.
(2) Assume that $\mathrm{char}(K)\neq 2$. Let $v$ be an arbitrary vertex of $E$. Then, it is clear that $v\in\mathbf{S}_{L_{K}(E)}$. Put $X_{0}=v$ and $X_{m+1}=X_{m}\circ v$, for all $m\geq 0$. Then, it can be checked that $X_{m}=2^{m}v$ for all $m\geq 0$. Since $\mathrm{char}(K)\neq 2$, it follows that $X_{m}\neq 0$ for all $m\geq 0$. Hence, $\mathbf{S}_{L_{K}(E)}$ is not Jordan nilpotent. ∎
To conclude this paper, we relate the solvability of the four structures $L_{K}(E)^{\circ}$, $\mathbf{S}_{L_{K}(E)}$, $L_{K}(E)^{-}$ and $\mathbf{K}_{L_{K}(E)}$.
Remark 3.
Theorem 4.11 is also true if we replace $\mathbf{S}_{L_{K}(E)}$ by the whole $L_{K}(E)^{\circ}$. This means $L_{K}(E)^{\circ}$ is Jordan solvable or Jordan nilpotent if and only if $\mathbf{S}_{L_{K}(E)}$ is. At the other extreme, in view of Theorem 4.6(2), if ${\rm char}(K)\neq 2$ then $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable but $L_{K}(E_{4})^{-}$ is not Lie solvable. These facts show that $\mathbf{K}_{L_{K}(E)}$ is Lie solvable for a wider class of graph $E$ than $L_{K}(E)^{-}$. Also, in some senses, the Lie algebra $\mathbf{S}_{L_{K}(E)}$ is bigger than the Jordan algebra $\mathbf{K}_{L_{K}(E)}$ in $L_{K}(E)$.
Funding. This research is funded by the Vietnam Ministry of Education and Training under grant number B2024-CTT-02.
Conflict of Interest. The authors have no conflict of interest to declare that are relevant to this article.
Data availability. The authors declare that the data supporting the findings of this study are available within the paper.

### v2

@@
 Every associative algebra $\mathcal{A}$ over a field $K$ can give rise to two new structures: the Lie algebra, denoted by $\mathcal{A}^{-}$, and the Jordan algebra, denoted by $\mathcal{A}^{\circ}$ over $K$.
-These structures can be obtained by defining new products within $\mathcal{A}$, namely the Lie bracket $[x,y]=xy-yx$ for Lie algebras and the Jordan product $x\circ y=xy+yx$ for Jordan algebras respectively, for every $x,y\in\mathcal{A}$.
+These structures can be obtained by defining new products within $\mathcal{A}$, namely the Lie bracket [x,y]=xyyx[x,y]=xy-yx for Lie algebras and the Jordan product $x\circ y=xy+yx$ for Jordan algebras respectively, for every $x,y\in\mathcal{A}$.
 A reasonable question arises: Are the properties of $\mathcal{A}$ as a ring mirrored in its Lie or Jordan algebra counterparts?
@@
 In a seminal paper [11] by Herstein, this question was answered affirmatively.
-Assume further that $\mathcal{A}$ is equipped with an involution $\star$; that is, a map ${}^{\star}:\mathcal{A}\to\mathcal{A}$ satisfying all below conditions:
+Assume further that $\mathcal{A}$ is equipped with an involution $\star$; that is, a map $^{\star}:\mathcal{A}\to\mathcal{A}$ satisfying the following conditions:
   1. (1)
@@
 $\mathbf{S}_{\mathcal{A}}=\{x\in\mathcal{A}:x^{\star}=x\}\text{ and }\mathbf{K}_{\mathcal{A}}=\{x\in\mathcal{A}:x^{\star}=-x\}$
-[…]
+be the sets of symmetric elements and skew-symmetric elements of $\mathcal{A}$ with respect to $\star$}, respectively.
+It is obvious that $\mathbf{K}_{\mathcal{A}}$ and $\mathbf{S}_{\mathcal{A}}$ are Lie subalgebra and Jordan subalgebra of $\mathcal{A}^{-}$ and $\mathcal{A}^{\circ}$ respectively.
+These are the most important structures of Lie or Jordan algebras arising from an associative algebra.
+In particular, when $\mathcal{A}$ is the full matrix ring $\mathbb{M}_{n}(K)$ and is the transpose operation, then $\mathbf{K}_{\mathcal{A}}$ is identified with the Lie algebra of skew-symmetric matrices which corresponds to the Lie group of orthogonal matrices.
+Under certain assumptions on $\mathcal{A}$, Lie and Jordan structures of $\mathbf{K}_{\mathcal{A}}$ and $\mathbf{S}_{\mathcal{A}}$ respectively were intensively investigated by Herstein in a series of papers ([7]-[11]) and by other authors in the references therein.
 In this paper, we consider the case when $\mathcal{A}$ is the Leavitt path algebra $L_{K}(E)$ of a directed graph $E$ over a field $K$.
+The concept of the Leavitt path algebra $L_{K}(E)$ associated with a graph $E$ and coefficients from a field $K$ made its debut in [2], sparking widespread interest among scholars.
+These algebras were concurrently and separately introduced in [6], despite being published two years following [2].
 Within the past few years, the Lie subalgebras of $L_{K}(E)^{-}$ arising from $L_{K}(E)$ have been investigated by many authors.
-Motivated by the groundbreaking work of Herstein [9], necessary and sufficient conditions for the Lie algebra $[L_{K}(E),L_{K}(E)]$ of $L_{K}(E)^{-}$ to be simple were given in [2] and [3], under the assumption that $E$ is row-finite.
+Motivated by the groundbreaking work of Herstein [11], necessary and sufficient conditions for the Lie algebra $L_{K}(E),L_{K}(E)]$ of $L_{K}(E)^{-}$ to be simple were given in [3] and [4], under the assumption that $E$ is row-finite.
 In another direction, the paper [12] completely identified the graph $E$ and the field $K$ for the Lie algebra $L_{K}(E)^{-}$ to be solvable.
@@
 A vertex $v$ is said to be regular if it is neither a sink nor an infinite emitter.
-A path $\mu$ in a graph $E$ is a finite sequence of edges $\mu=e_{1}e_{2}\dots e_{n}$ with $r(e_{i})=s(e_{i+1})$ for all $11\leq i\leq n-$ 1.
+A path $\mu$ in a graph $E$ is a finite sequence of edges $\mu=e_{1}e_{2}\dots e_{n}$ with $r(e_{i})=s(e_{i+1})$ for all $1\leq i\leq n-1$.
 We set $s(\mu):=s(e_{1})$ and $r(\mu):=r(e_{n})$.
@@
 2.2. Lie and Jordan structures of associative algebras
-Every associative algebra $\mathcal{A}$ over a field $K$ is endowed with a Lie algebra structure for the Lie bracket $[a,b]=ab-ba$ for all $a,b\in\mathcal{A}$.
+Every associative algebra $\mathcal{A}$ over a field $K$ is endowed with a Lie algebra structure for the Lie bracket [a,b]=abba[a,b]=ab-ba for all $a,b\in\mathcal{A}$.
 We denote this Lie algebra by $\mathcal{A}^{-}$.
-For $K$-subspaces $S$ and $T$ of $\mathcal{A}$, we defined $[S,T]={\rm span}_{K}\{ab-ba:a\in S,b\in T\}$.
-A $K$-subspace $\mathcal{L}$ of $\mathcal{A}$ is called a Lie subalgebra of $\mathcal{A}^{-}$ if $[a,b]\in\mathcal{L}$ for all $a,b\in\mathcal{L}$.
+For $K$-subspaces $S$ and $T$ of $\mathcal{A}$, we define $S,T]={\rm span}_{K}\{ab-ba:a\in S,b\in T\}$.
+A $K$-subspace $\mathcal{L}$ of $\mathcal{A}$ is called a Lie subalgebra of $\mathcal{A}^{-}$ if $a,b]\in\mathcal{L}$ for all $a,b\in\mathcal{L}$.
 For the Lie subalgebra $\mathcal{L}$ of $\mathcal{A}^{-}$, we define inductively the two series
@@
 The Lie subalgebra $\mathcal{L}$ is called Lie solvable (resp., Lie nilpotent) if there exists $n\geq 0$ (resp., $m\geq 0$) such that $\mathcal{L}^{(n)}=0$ (resp., $\mathcal{L}^{m}=0$).
-If such the smallest $n$ (resp., $m$) exists, we will say that $\mathcal{L}$ is Lie solvable of index $n$ (resp., Lie nilpotent of index $m$).
+If such a smallest $n$ (resp., $m$) exists, we will say that $\mathcal{L}$ is Lie solvable of index $n$ (resp., Lie nilpotent of index $m$).
 Similarly, a Jordan product can be defined on $\mathcal{A}$ by $a\circ b=ab+ba$, for all $a,b\in\mathcal{A}$.
 If $\mathcal{A}$ is viewed as a Jordan algebra for the product $\circ$, we denote it by $\mathcal{A}^{\circ}$.
-Note that in the case ${\rm char}(K)=2$, the two structures $\mathcal{A}^{\circ}$ and $\mathcal{A}^{-}$ coincides.
+Note that in the case ${\rm char}(K)=2$, the two structures $\mathcal{A}^{\circ}$ and $\mathcal{A}^{-}$ coincide.
 A $K$-subspace $\mathcal{J}$ of $\mathcal{A}$ is called a Jordan subalgebra of $\mathcal{A}^{\circ}$ if $a\circ b\in\mathcal{J}$ for all $a,b\in\mathcal{J}$.
-For two $K$-subspaces $S$ and $T$ of $\mathcal{A}$, we defined
+For two $K$-subspaces $S$ and $T$ of $\mathcal{A}$, we define
 $S\circ T={\rm span}_{K}\{ab+ba:a\in S,b\in T\}.$
@@
 A homomorphism of $K$-algebras with involution $\varphi:(\mathcal{A},^{\star})\to(\mathcal{B},^{\natural})$ is a $K$-algebra homomorphism $\varphi:\mathcal{A}\to\mathcal{B}$ such that $(\varphi(ab))^{\natural}=\varphi(b^{\star})\varphi(a^{\star})$.
-It is obvious that if $(\mathcal{A},^{\star})$ is isomorphic to $(\mathcal{B},^{\natural})$ as $K$-algebras with involution, then the Lie subalgebras $\mathbf{K}_{\mathcal{A}}$ and $\mathbf{K}_{\mathcal{B}}$, with the respect involutions and $\natural$ respectively, are also isomorphic as Lie algebras over $K$.
-Similarly, the Jordan subalgebras $\mathbf{S}_{\mathcal{A}}$ and $\mathbf{S}_{\mathcal{B}}$ of $\mathcal{A}^{\circ}$ and $\mathcal{B}^{\circ}$, with the respect involutions and $\natural$ respectively, are also isomorphic as Jordan algebras over $K$.
+It is obvious that if $(\mathcal{A},^{\star})$ is isomorphic to $(\mathcal{B},^{\natural})$ as $K$-algebras with involution, then the Lie subalgebras $\mathbf{K}_{\mathcal{A}}$ and $\mathbf{K}_{\mathcal{B}}$, with respect to the involutions and $\natural$ respectively, are also isomorphic as Lie algebras over $K$.
+Similarly, the Jordan subalgebras $\mathbf{S}_{\mathcal{A}}$ and $\mathbf{S}_{\mathcal{B}}$ of $\mathcal{A}^{\circ}$ and $\mathcal{B}^{\circ}$, with respect to the involutions and $\natural$ respectively, are also isomorphic as Jordan algebras over $K$.
 3. Lie Solvability and Jordan Solvability of $\mathbb{M}_{n}(\mathcal{A})$
@@
 Let $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ be the Lie subalgebra of $\mathbb{M}_{n}(\mathcal{A})^{-}$ with respect to the involution $\natural$ on $\mathbb{M}_{n}(\mathcal{A})$ as defined in (3.1).
-In this section, we gives conditions on $n$ and $\mathcal{A}$ for $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ to be Lie solvable.
-First, we fix the following notation: In $\mathbb{M}_{n}(\mathcal{A})$, we denote by $E_{ij}$ the matrix whose $(i,j)$-entry is $1$ and the other entries are 00.
+In this section, we give conditions on $n$ and $\mathcal{A}$ for $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ to be Lie solvable.
+First, we fix the following notation: In $\mathbb{M}_{n}(\mathcal{A})$, we denote by $E_{ij}$ the matrix whose $(i,j)$-entry is $1$ and the other entries are 0.
 The following proposition is the main result of this section:
@@
 Let $\mathbf{K}_{\mathbb{M}_{n}(\mathcal{A})}$ be as defined above.
-Then, the following assertions holds:
+Then, the following assertions hold:
   1. (a)
@@
     Assume that ${\rm char}(K)\neq 2$.
-If the involution on $\mathcal{A}$ is trivial, then $[\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})},\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}]=0$; while if the involution on $\mathcal{A}$ is non-trivial, then $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is not Lie solvable.
+If the involution on $\mathcal{A}$ is trivial, then $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})},\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}]=0$; while if the involution on $\mathcal{A}$ is non-trivial, then $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is not Lie solvable.
 Proof.
@@
 It follows that
-(3.4) $[X_{1},X_{2}]=\left(\begin{array}[]{cc}s_{2}s_{1}^{\natural}-s_{1}s_{2}^{\natural}&0\\[8.5359pt] 2(r_{1}s_{2}^{\natural}-s_{1}^{\natural}r)&s_{2}^{\natural}s_{1}-s_{1}^{\natural}s_{2}\end{array}\right),\text{ where }$
+(3.4) $X_{1},X_{2}]=\left(\begin{array}[]{cc}s_{2}s_{1}^{\natural}-s_{1}s_{2}^{\natural}&0\\[8.5359pt] 2(r_{1}s_{2}^{\natural}-s_{1}^{\natural}r)&s_{2}^{\natural}s_{1}-s_{1}^{\natural}s_{2}\end{array}\right),\text{ where }$
 (3.5) $s_{2}s_{1}^{\natural}-s_{1}s_{2}^{\natural}\hskip 2.84544pt=(v_{2}v_{1}^{\natural}-v_{1}v_{2}^{\natural})(a_{1}-c_{1})(c_{2}-a_{2})$
@@
 (b) and (c): Assume that ${\rm char}(K)=2$.
-As $02(r_{1}s_{2}^{\natural}-s_{1}^{\natural}r)=$ 0, it follows from (3.4) that $[X_{1},X_{2}]$ is a diagonal matrix.
-Similarly, we also obtain that $[X_{3},X_{4}]$ is also a diagonal matrix.
-Thus $[[X_{1},X_{2}],[X_{3},X_{4}]]=0$, which means that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(3)}=0$.
+As $2(r_{1}s_{2}^{\natural}-s_{1}^{\natural}r)=0$, it follows from (3.4) that $X_{1},X_{2}]$ is a diagonal matrix.
+Similarly, we also obtain that $X_{3},X_{4}]$ is also a diagonal matrix.
+Thus $X_{1},X_{2}],[X_{3},X_{4}]]=0$, which means that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(3)}=0$.
 We divide our situation into two possible cases:
@@
 As $b_{i}^{\natural}=b_{i}$ and $v_{i}^{\natural}=v_{i}$, we get from (3.3) that $r_{i}=0$.
-It follows from (3.2) that $X_{i}=\left(\begin{array}[]{cc}0&s_{i}\\ -s_{i}^{\natural}&0\end{array}\right)$, and so $[X_{1},X_{2}]=0$; or equivalently, we have $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(2)}=0$.
+It follows from (3.2) that $X_{i}=\left(\begin{array}[]{cc}0&s_{i}\\ -s_{i}^{\natural}&0\end{array}\right)$, and so $X_{1},X_{2}]=0$; or equivalently, we have $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}^{(2)}=0$.
 In other words, $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is Lie solvable of index 2.
@@
 ${\rm char}(K)=2$ and the involution on $\mathcal{A}$ is non-trivial.
-In this case, suitable choices of the matrices $A_{1},A_{2},B_{1},B_{2}$ yield that $[X_{1},X_{2}]\neq 0$.
+In this case, suitable choices of the matrices $A_{1},A_{2},B_{1},B_{2}$ yield that $X_{1},X_{2}]\neq 0$.
 For example, if we choose $a_{1}=c_{1}$, $a_{2}=c_{2}$, $b_{1}=1$, $u_{1}\neq w_{1}$, $w_{1}=w_{2}=0$, $0\neq u_{1}=u_{2}\in K$, $b_{2}\neq b_{2}^{\natural}$, then it follows from (3.5) that $s_{2}s_{1}^{\natural}-s_{1}s_{2}^{\natural}=(b_{2}^{\natural}-b_{2})u_{1}^{2}\neq 0$.
-Therefore, we get from (3.4) that $[X_{1},X_{2}]\neq 0$.
+Therefore, we get from (3.4) that $X_{1},X_{2}]\neq 0$.
 It follows that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}$ is Lie solvable of index 3.
@@
 ${\rm char}(K)\neq 2$ and the involution on $\mathcal{A}$ is trivial.
-Because $a^{\natural}=a$, the equation $a=-a^{\natural}$ is equivalent to $02a=$ 0, and hence $a=0$.
+Because $a^{\natural}=a$, the equation $a=-a^{\natural}$ is equivalent to $2a=0$, and hence $a=0$.
 Similarly, we also obtain that $c=0$.
@@
 $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}=\left\{\begin{pmatrix}0&b\\ -b&0\end{pmatrix}\;|\;b\in\mathcal{A}\right\}.$
-It follows that $[\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})},\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}]=0$.
+It follows that $\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})},\mathbf{K}_{\mathbb{M}_{2}(\mathcal{A})}]=0$.
 Case 4.
@@
     Suppose that $n=2$.
-If ${\rm char}(K)\neq 2$, then $[\mathbf{K}_{\mathbb{M}_{n}(K)},\mathbf{K}_{\mathbb{M}_{n}(K)}]=0$; and, if ${\rm char}(K)=2$, then $\mathbf{K}_{\mathbb{M}_{n}(K)}$ is Lie solvable of index $2$ but not Lie nilpotent.
+If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{\mathbb{M}_{n}(K)},\mathbf{K}_{\mathbb{M}_{n}(K)}]=0$; and, if ${\rm char}(K)=2$, then $\mathbf{K}_{\mathbb{M}_{n}(K)}$ is Lie solvable of index $2$ but not Lie nilpotent.
   2. (2)
@@
 Let $K[x,x^{-1}]$ be the Laurent polynomial $K$-algebra.
-Let $\mathbf{K}_{\mathbb{M}_{n}(K[x,x^{-1}])}$ be the Lie subalgebra of skew-symmetric elements of $\mathbb{M}_{n}(K[x,x^{-1}])^{-}$ with respect to the involution defined by ((((3.7)))).
+Let $\mathbf{K}_{\mathbb{M}_{n}(K[x,x^{-1}])}$ be the Lie subalgebra of skew-symmetric elements of $\mathbb{M}_{n}(K[x,x^{-1}])^{-}$ with respect to the involution defined by ((3.7)).
 Then, the following assertions hold:
@@
 4. Lie Solvability and Jordan Solvability of $L_{K}(E)$
-Throughout of this section, we will denote by $\mathbf{K}_{L_{K}(E)}$ and $\mathbf{S}_{L_{K}(E)}$ the Lie subalgebra of $L_{K}(E)^{-}$ and the Jordan subalgebra of $L_{K}(E)^{\circ}$with respect to standard involution on $L_{K}(E)$ as defined before, respectively.
+Throughout this section, we will denote by $\mathbf{K}_{L_{K}(E)}$ and $\mathbf{S}_{L_{K}(E)}$ the Lie subalgebra of $L_{K}(E)^{-}$ and the Jordan subalgebra of $L_{K}(E)^{\circ}$with respect to standard involution on $L_{K}(E)$ as defined before, respectively.
 The main aim of this section is to provide a necessary and sufficient condition on the graph $E$ and the field $K$ for $\mathbf{K}_{L_{K}(E)}$ and $\mathbf{S}_{L_{K}(E)}$ to be Lie and Jordan solvable, respectively.
-We start this section a lemma which determines a $K$-basis of $L_{K}(E)$:
+We start this section with a lemma which determines a $K$-base of $L_{K}(E)$:
 Lemma 4.1 ([1, Corollary 1.5.12]).
@@
 For each $v\in{\rm Reg}(E)$, let $\{e_{1}^{v},\dots,e_{n_{v}}^{v}\}$ be an enumeration of the elements of $s^{-1}(v)$.
-Then, a $K$-basic of $L_{K}(E)$ is given by the family
+Then, a $K$-basis of $L_{K}(E)$ is given by the family
 $\mathscr{B}=\mathscr{A}\backslash\{\lambda e_{n_{v}}^{v}(e_{n_{v}}^{v})^{*}\nu^{*}\;:\;r(\lambda)=r(\nu)=v\in{\rm Reg}(E)\}.$
 The next lemma provides a necessary condition for $\mathbf{K}_{L_{K}(E)}$ to be Lie solvable.
-The ideas used to prove of this lemma come from the proof of [12, Lemma 2.2 and Theorem 2.3].
+The ideas used to prove this lemma come from the proof of [12, Lemma 2.2 and Theorem 2.3].
 It can be said that, in $L_{K}(E)$, we can easily choose a suitable subalgebra which is isomorphic to the matrix ring $\mathbb{M}_{3}(K)$.
@@
 Lemma 4.2.
-. Let $L_{K}(E)$ be the Leavitt path algebra of a graph $E$ with coefficients in a field $K$. Then, $\mathbf{K}_{L_{K}(E)}$ is not Lie solvable if one of the following cases occurs:
+Let $L_{K}(E)$ be the Leavitt path algebra of a graph $E$ with coefficients in a field $K$. Then, $\mathbf{K}_{L_{K}(E)}$ is not Lie solvable if one of the following cases occurs:
   1. (1)
@@
     $E$ contains one of following subgraphs:
-[…]
-[…]
-[…]
+    $F_{1}:\bullet\bullet\bulletvveeff$
+    $F_{2}:\bullet\bullet\bulletvveeff$
+    $F_{3}:\bullet\bulletvveeff$
 Proof.
@@
 Then, $v,f,ef$ are all paths ending at $v$.
-It is straightforward to check the subalgebra $B$ of $L_{K}(E)$ generated by $v$, $f$, $ef$, $f^{*}$, $f^{*}e^{*}$, $ff^{*}$, $eff^{*}$, $ff^{*}e^{*}$, $eff^{*}e^{*}$ is isomorphic to $\mathbb{M}_{3}(K)$ via map: $eff^{*}e^{*}\mapsto E_{11}$, $ff^{*}\mapsto E_{22}$, $v\mapsto E_{33}$, $eff^{*}\mapsto E_{12}$, $ff^{*}e^{*}\mapsto E_{21}$, $f\mapsto E_{23}$, $f^{*}\mapsto E_{32}$, $ef\mapsto E_{13}$, and $f^{*}e^{*}\mapsto E_{31}$ (see also a similar techniques used in the proof of [1, Lemma 2.6.4]).
+It is straightforward to check the subalgebra $B$ of $L_{K}(E)$ generated by $v$, $f$, efef, $f^{*}$, $^{*}e^{*}, ff$$f^{*}$, $^{*}, ffeff^{*}e^{*}$, $eff^{*}e^{*}$ is isomorphic to $\mathbb{M}_{3}(K)$ via map: $eff^{*}e^{*}\mapsto E_{11}$, $ff^{*}\mapsto E_{22}$, $v\mapsto E_{33}$, $eff^{*}\mapsto E_{12}$, $ff^{*}e^{*}\mapsto E_{21}$, $f\mapsto E_{23}$, $f^{*}\mapsto E_{32}$, $ef\mapsto E_{13}$, and $f^{*}e^{*}\mapsto E_{31}$ (see also a similar techniques used in the proof of [1, Lemma 2.6.4]).
 The restriction of the involution to $B$ is also an involution on $B$ for which $\mathbf{K}_{B}\subseteq\mathbf{K}_{L_{K}(E)}$.
@@
 $E_{2}:\bullet$
-[…]
+$E_{3}:\bullet\bullet$
 […]
@@
 Lemma 4.3.
-Let $\mathcal{A}$ be a $K$-algebra with an involution , and $\mathbf{J}$ be an indexing set of arbitrary cardinality.
+Let $\mathcal{A}$ be a $K$-algebra with an involution $\star$, and $\mathbf{J}$ be an indexing set of arbitrary cardinality.
 Let $I_{j},j\in\mathbf{J}$, be ideals of $\mathcal{A}$ satisfying the property that for each $j$, if $a\in I_{j}$ then $a^{\star}\in I_{j}$.
@@
 If $M=\bigoplus_{j\in\mathbf{J}}I_{j}$, then $\mathbf{K}_{M}=\bigoplus_{j\in\mathbf{J}}\mathbf{K}_{I_{j}}$.
-Consequently, $\mathbf{K}_{M}$ is Lie solvable ((((nilpotent)))) if and only if $\mathbf{K}_{I_{j}}$ is Lie solvable ((((nilpotent)))) of a common index for all $j\in\mathbf{J}$.
+Consequently, $\mathbf{K}_{M}$ is Lie solvable ((nilpotent)) if and only if $\mathbf{K}_{I_{j}}$ is Lie solvable ((nilpotent)) of a common index for all $j\in\mathbf{J}$.
 Proof.
@@
   2. (2)
-    $\mathbf{K}_{L_{K}(E_{3})}$ is Lie solvable ((((of index $3$)))) if and only if $\mathrm{char}(K)=2$. Additionally, $\mathbf{K}_{L_{K}(E_{3})}$ is not Lie nilpotent.
+    $\mathbf{K}_{L_{K}(E_{3})}$ is Lie solvable ((of index 33)) if and only if $\mathrm{char}(K)=2$. Additionally, $\mathbf{K}_{L_{K}(E_{3})}$ is not Lie nilpotent.
 Proof.
@@
     If $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable, then it is Lie solvable of index $l\leq 3$.
-More precisely, if $|I|$ is finite and ${\rm char}(K)\neq 2$ ((((resp., ${\rm char}(K)=2$)))), then $l=1$ ((((resp., $l=2$)))).
-If $|I|$ is infinite and ${\rm char}(K)\neq 2$ ((((reps., ${\rm char}(K)=2$)))), then $l=2$ ((((resp., $l=3$)))).
-Additionally, $\mathbf{K}_{L_{K}(E_{4})}$ is Lie nilpotent if and only if $|I|$ is finite and ${\rm char}(K)\neq 2$, if and only if $[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]=1$.
+More precisely, if $|I|$ is finite and ${\rm char}(K)\neq 2$ ((resp., ${\rm char}(K)=2$)), then $l=1$ ((resp., $l=2$)).
+If $|I|$ is infinite and ${\rm char}(K)\neq 2$ ((reps., ${\rm char}(K)=2$)), then $l=2$ ((resp., $l=3$)).
+Additionally, $\mathbf{K}_{L_{K}(E_{4})}$ is Lie nilpotent if and only if $|I|$ is finite and ${\rm char}(K)\neq 2$, if and only if $\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]=1$.
   2. (2)
-    $\mathbf{K}_{L_{K}(E_{5})}$ is Lie solvable ((((of index $l\leq 4$)))) if and only if ${\rm char}(K)=2$.
-More precisely, if $|J|$ is finite ((((resp., infinite)))), then $l=3$ ((((resp., $l=4$)))).
+    $\mathbf{K}_{L_{K}(E_{5})}$ is Lie solvable ((of index $l\leq 4$)) if and only if ${\rm char}(K)=2$.
+More precisely, if $|J|$ is finite ((resp., infinite)), then $l=3$ ((resp., $l=4$)).
 Additionally, $\mathbf{K}_{L_{K}(E_{5})}$ is not Lie nilpotent.
   3. (3)
-    $\mathbf{K}_{L_{K}(E_{6})}$ is Lie solvable ((((of index $l\leq 4$)))) if and only if ${\rm char}(K)=2$.
-More precisely, if both $|I|$ and $|J|$ are finite ((((resp., either $|I|$ or $|J|$ is infinite)))), then $l=3$ ((((resp., $l=4$)))).
+    $\mathbf{K}_{L_{K}(E_{6})}$ is Lie solvable ((of index $l\leq 4$)) if and only if ${\rm char}(K)=2$.
+More precisely, if both $|I|$ and $|J|$ are finite ((resp., either $|I|$ or $|J|$ is infinite)), then $l=3$ ((resp., $l=4$)).
 Additionally, $\mathbf{K}_{L_{K}(E_{6})}$ is not Lie nilpotent.
@@
 Then, according to Lemma 4.3 and Corollary 3.2, we conclude that $\mathbf{K}_{M}=\bigoplus_{i=1}^{n}\mathbf{K}_{I_{i}}$, and that $\mathbf{K}_{M}$ is Lie solvable.
-Moreover, in view of Corollary 3.2, if ${\rm char}(K)\neq 2$, then $[\mathbf{K}_{M},\mathbf{K}_{M}]=0$; and, if ${\rm char}(K)=2$ then $\mathbf{K}_{M}$ is Lie solvable of index $2$ but not Lie nilpotent.
+Moreover, in view of Corollary 3.2, if ${\rm char}(K)\neq 2$, then $\mathbf{K}_{M},\mathbf{K}_{M}]=0$; and, if ${\rm char}(K)=2$ then $\mathbf{K}_{M}$ is Lie solvable of index $2$ but not Lie nilpotent.
 Now, we consider two possible cases:
@@
 Case 2. $n$ is infinite. In this case, $u$ is an infinite emitter and $u\not\in M$. This implies that $L_{K}(E_{4})=Ku\oplus M$ as $K$-vector spaces. It follows that
-$[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]\subseteq[L_{K}(E_{4}),L_{K}(E_{4})]=[M,M]\subseteq M.$
-As $M$ is Lie solvable, we conclude that $[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]$, and hence $\mathbf{K}_{L_{K}(E_{4})}$, is Lie solvable.
+$\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]\subseteq[L_{K}(E_{4}),L_{K}(E_{4})]=[M,M]\subseteq M.$
+As $M$ is Lie solvable, we conclude that $\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]$, and hence $\mathbf{K}_{L_{K}(E_{4})}$, is Lie solvable.
 To determine the solvable index of $\mathbf{K}_{L_{K}(E_{4})}$, we consider the following subcases:
@@
 ${\rm char}(K)=2$.
-As $[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]\subseteq\mathbf{K}_{M}$ which is Lie solvable of index $2$, it follows that $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable of index at most $3$.
+As $\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]\subseteq\mathbf{K}_{M}$ which is Lie solvable of index $2$, it follows that $\mathbf{K}_{L_{K}(E_{4})}$ is Lie solvable of index at most $3$.
 Because $v\in\mathbf{K}_{L_{K}(E_{4})}\backslash\mathbf{K}_{M}$, we conclude that $\mathbf{K}_{L_{K}(E_{4})}\neq[\mathbf{K}_{L_{K}(E_{4})},\mathbf{K}_{L_{K}(E_{4})}]$.
@@
   1. (1)
-    If ${\rm char}(K)=2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((((of index $\leq 4$)))) if and only if $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$.
+    If ${\rm char}(K)=2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((of index $\leq 4$)) if and only if $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$.
 Specially, the index is $4$ precisely when $E$ contains either $E_{5}$ or $E_{6}$ with infinite $|I|$ or $|J|$.
   2. (2)
-    If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((((of index $\leq 2$)))) if and only if $E$ is a disjoint union of graphs $E_{1}$, $E_{2}$ and $E_{4}$.
+    If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((of index $\leq 2$)) if and only if $E$ is a disjoint union of graphs $E_{1}$, $E_{2}$ and $E_{4}$.
 Specially, the index is $2$ precisely when $E$ contains $E_{4}$ with infinite $|I|$.
@@
   1. (1)
-    If ${\rm char}(K)=2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((((of index $\leq 3$)))) if and only if $E$ is a disjoint union graphs $E_{1},\dots,E_{6}$ with finite $|I|$ and $|J|$.
+    If ${\rm char}(K)=2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((of index $\leq 3$)) if and only if $E$ is a disjoint union graphs $E_{1},\dots,E_{6}$ with finite $|I|$ and $|J|$.
 In this case,
     $L_{K}(E)\cong\bigoplus_{I_{1}}K\oplus\bigoplus_{I_{2}}K[x,x^{-1}]\oplus\bigoplus_{I_{3}}\mathbb{M}_{2}(K)\oplus\bigoplus_{I_{4}}\mathbb{M}_{2}(K[x,x^{-1}]),$
-    for suitable ((((possibly infinite)))) indexing sets $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$. Moreover, the index is $3$ precisely when $E$ contains $E_{3}$, $E_{5}$ or $E_{6}$.
+    for suitable ((possibly infinite)) indexing sets $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$. Moreover, the index is $3$ precisely when $E$ contains $E_{3}$, $E_{5}$ or $E_{6}$.
   2. (2)
-    If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((((of index $\leq 2$)))) if and only if $E$ is a disjoint union of graphs $E_{1}$, $E_{2}$ and $E_{4}$ with finite $|I|$.
+    If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie solvable ((of index $\leq 2$)) if and only if $E$ is a disjoint union of graphs $E_{1}$, $E_{2}$ and $E_{4}$ with finite $|I|$.
 In this case, we have
     $L_{K}(E)\cong\bigoplus_{J_{1}}K\oplus\bigoplus_{J_{2}}K[x,x^{-1}]\oplus\bigoplus_{J_{3}}\mathbb{M}_{2}(K),$
-    for suitable ((((possibly infinite)))) indexing sets $J_{1}$, $J_{2}$, $J_{3}$. Moreover, the index is $2$ precisely when $E$ contains $E_{3}$.
+    for suitable ((possibly infinite)) indexing sets $J_{1}$, $J_{2}$, $J_{3}$. Moreover, the index is $2$ precisely when $E$ contains $E_{3}$.
 The following theorem provides a necessary and sufficient condition for $\mathbf{K}_{L_{K}(E)}$ to be Lie nilpotent.
@@
     If ${\rm char}(K)\neq 2$, then $\mathbf{K}_{L_{K}(E)}$ is Lie nilpotent if and only if $E$ is a disjoint union graphs $E_{1}$, $E_{2}$, and $E_{4}$ with finite $n$.
-In particular, $\mathbf{K}_{L_{K}(E)}$ is Lie nilpotent if and only if $[\mathbf{K}_{L_{K}(E)},\mathbf{K}_{L_{K}(E)}]=0$.
+In particular, $\mathbf{K}_{L_{K}(E)}$ is Lie nilpotent if and only if $\mathbf{K}_{L_{K}(E)},\mathbf{K}_{L_{K}(E)}]=0$.
 Corollary 4.9.
@@
   1. (1)
-    If ${\rm char}(K)=2$, then $\mathbf{S}_{L_{K}(E)}$ is Jordan solvable ((((resp., nilpotent)))) if and only if $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$ ((((resp., $E_{1}$ and $E_{2}$)))).
+    If ${\rm char}(K)=2$, then $\mathbf{S}_{L_{K}(E)}$ is Jordan solvable ((resp., nilpotent)) if and only if $E$ is a disjoint union of graphs $E_{1},\dots,E_{6}$ ((resp., $E_{1}$ and $E_{2}$)).
   2. (2)
@@
 Also, in some senses, the Lie algebra $\mathbf{S}_{L_{K}(E)}$ is bigger than the Jordan algebra $\mathbf{K}_{L_{K}(E)}$ in $L_{K}(E)$.
-Funding. This research is funded by the Vietnam Ministry of Education and Training under grant number B2024-CTT-02.
-Conflict of Interest. The authors have no conflict of interest to declare that are relevant to this article.
-Data availability. The authors declare that the data supporting the findings of this study are available within the paper.
+Funding. This work is funded by Vietnam National Foundation for Science and Technology Development (NAFOSTED) under Grant No. 101.04-2025.41.